Vino Veritas Company, a U.S.-based importer of wines and spirits, placed an order with a French supplier for 1,500 cases of wine at a price of 250 euros per case. The total purchase price is 375,000 euros. Relevant exchange rates for the euro are as follows:

Spot rate Forward rate to October 31 Call Option preiumum for october 31 strick price 1.25
9/15/2015 1.25 1.31 0.040
9/30/2013 1.30 1.34 0.075
10/31/2015 1.35 1.35 0.100

9/30

Forward Contract

Gain on forward contract

10/31

Forward contract

Gain on forward contract

Please she work.
Answer #1

Answer:

A US based Importer imports goods from a French supplier and ends in payment of euros.

If he ends his payment on 9/30/2015:

Spot rate would come to 1.30 $/€

Alternative 1:
Forward rate is 1.34$/€
In forward contract his payment would be = € 375,000 * 1.34$/€ = $ 502,500

Alternative 2:

In options :

strike price is 1.25$/€ and option premium is 0.075$

Since strike price is lesser than spot price, the call option is beneficial.

payment under call option is = € 375,000 * (1.25$/€ + 0.075$) = $496,875

Alternative 3:

If no forward cover or option cover is taken,

His payment would be in spot rate= € 375,000 * 1.30$/€ = $ 487,500

Conclusion:
From above analysis if forward cover is taken the importer would end up in paying higher amount than compared to other alternatives.

Loss on forward contract would be = $ 487,500 - $502,500 = ($ 15,000)

(on comparison to best alternative)


If he ends his payment on 10/31/2015:

There would not be any gain or loss on entering into forward contract because on this date all the alternatives have same price for the purchase of € or selling a $.

{ spot rate = 1.35 $/€

Forward rate =  1.35 $/€

option contract = strike price + premium =  1.25 $/€ + 0.100$ =  1.35 $/€}

Note: It is assumed that by any mistake in the question provided the date is 9/30/2015 in second row instead of 9/30/2013.